OpenStudy (anonymous):

Please help me... Solve the following equation for the value of x: (ln x)^2 = ln x^2

OpenStudy (anonymous):

\[\left( \ln x \right)^{2}=\ln x ^{2}\] just to better understand this is what the equation really looks like

OpenStudy (amoodarya):